mhchen:

Solve x^2 + 4x - 10 = 0 please @pooja195

pooja195:

Hi there ;) Welcome to Question Cove Any ideas as to where we should start?

pooja195:

Hmm factoring is not going to work here, we will need to use the quadratic formula which looks like this: \[\huge~\rm~x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] \(\normalsize\color{ slate }{{\bbox[5pt, lightcyan ,border:2px solid black ]{ \LARGE{x=~} \huge{ \frac{-\color{magenta}{b} \pm\sqrt{ \color{magenta}{~b} ^2-4 \color{blue}{a} \color{red}{c}~}}{2 \color{blue}{a}} }~ }}}\) when the equation is \(\large\color{black}{ \color{blue}{a} x^2+ \color{magenta}{b}x+ \color{red}{c}=0 }\). Are you familiar with this?

mhchen:

Hot dayum, lol. Yes I am familiar with that.

pooja195:

Can you tell me the a,b,c values? :P

mhchen:

a is 1, b is 4, c is -10

pooja195:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @pooja195 Hmm factoring is not going to work here, we will need to use the quadratic formula which looks like this: \[\huge~\rm~x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] \(\normalsize\color{ slate }{{\bbox[5pt, lightcyan ,border:2px solid black ]{ \LARGE{x=~} \huge{ \frac{-\color{magenta}{4} \pm\sqrt{ \color{magenta}{~4} ^2-4 \color{blue}{(1)} \color{red}{(-10)}~}}{2 \color{blue}{(1)}} }~ }}}\) when the equation is \(\large\color{black}{ \color{blue}{a} x^2+ \color{magenta}{b}x+ \color{red}{c}=0 }\). Are you familiar with this? \(\color{#0cbb34}{\text{End of Quote}}\) Plug it in B)

mhchen:

Wow, I got my answer of \[-2\pm \sqrt{14}\] Thanks Pooja.

pooja195:

yw ;) (;